# Nodal superconducting gap structure and topological surface states of UTe2

### v1

@@ III.2 Superconducting topological surface states
As a result of the odd parity of the superconducting order parameter, the boundaries of this topological system may exhibit zero-energy bound states in the form of Majorana-Dirac surface cones, that is, states with linear dispersion that form around time-reversal invariant momenta (TRIM) in the surface BZ [48, 2, 25, 49, 28, 27].
@@
[…]
In order to gain insight into the topological nature of the surface bound states and determine whether there are any qualitative difference between the B2u versus B3u phases, we consider the symmetry representations for the two different phases. We will exploit that the (0-11) surface preserves only the $\mathcal{M}_{x}$ crystal symmetry along with time-reversal $\mathcal{T}$ and the particle-hole anti-symmetry $\mathcal{P}$. Since the symmetry properties of the B2u and B3u phases are different under $\mathcal{M}_{x}$, this leads to qualitative differences that are robust against microscopic changes of the model that preserve $\mathcal{M}_{x}$. Their distinct symmetries under $\mathcal{M}_{x}$ can be seen directly from Fig. 4 with B2u (B3u) being odd (even) under $\mathcal{M}_{x}$. Specifically, we will see that the B2u phase features Majorana flat bands around the $\Gamma$ point, while the B3u exhibits a Majorana-Dirac surface cone with linear dispersion in both $k_{x}$ and $k_{c^{*}}$.
First, we consider the algebra of the representations of the symmetries in the bulk. Following Ref. [49], we express the commutation relation between the BdG representation $U$ of the crystal symmetry $g$ and the particle-hole anti-symmetry $\mathcal{P}$ in terms of the character $\chi(g)$ as
$U(g)U(\mathcal{P})=\chi(g)U(\mathcal{P})\mathcal{U}(g),$ (17)
where the representation of particle-hole is given by $U(\mathcal{P})=\tau_{1}K$, where $\tau_{1}$ is the first Pauli matrices in particle-hole space and $K$ is complex conjugation. Due to the presence of strong SOC, the spin and momenta transform together, resulting in the normal state representation of the mirror symmetries given by $u(\mathcal{M}_{i})=i\sigma_{i}$, where $\sigma_{i}$ are the Pauli matrices in spin-space. Both the bulk and the surface respect time-reversal symmetry and particle-hole anti-symmetry and therefore also the chiral anti-symmetry defined as $\mathcal{C}=\mathcal{T}\mathcal{P}$ [50]. Note that chiral symmetry leaves the momentum invariant, acting only on the internal degrees of freedom. Since time-reversal symmetry $U(\mathcal{T})=i\sigma_{y}K$ commutes with $u(\mathcal{M}_{x})=i\sigma_{x}$, the commutation relation between $\mathcal{C}$ and $\mathcal{M}_{x}$ are given by
@@
where $\chi(\mathcal{M}_{x})=1$ for B2u and $\chi(\mathcal{M}_{x})=-1$ for B3u as seen from Tab. 5. In the basis of the Majorana zero modes, a similar relation must hold for the representation of the symmetries of the surface, which in turn constrains the dispersion of the surface states. To see this, we write an effective surface Hamiltonian $H_{s}(\mathbf{k^{\parallel}})$ around the $\Gamma$ point. In both the B2u and B3u cases, we observe two localized surface states and we can therefore write it as a two-band model in terms of the Pauli matrices $\eta_{i}$. We denote the symmetry representation on the surface states $\mathcal{O}(g)$ where $g=\mathcal{M}_{x},\mathcal{P},\mathcal{T}$ are the set of symmetry operations preserved by the (0-11) cleave plane. We define the basis such that time-reversal and particle-hole representations are given by
$\mathcal{O}(\mathcal{P})=\eta_{0}K,$
@@
The above argument relies on the fact that only one crystal symmetry remains on the surface, under which B3u and B2u transform differently. This should be contrasted to a related discussion in Ref. [28] where the authors calculate the surface states on the (001) surface. That is, a surface preserving both $\mathcal{M}_{x}$ and $\mathcal{M}_{y}$ symmetries, such that no qualitative difference can be made between the B2u and B3u phases. Indeed, they find that both phases generically host flat bands around the $\Gamma$ point [28].
@@ Acknowledgements.
We acknowledge fruitful discussions with Daniel Agterberg, Séamus Davis, Peter J. Hirschfeld, and Shuqui Wang. H. C. acknowledges support from the Novo Nordisk Foundation grant NNF20OC0060019. A.K. acknowledges support by the Danish National Committee for Research Infrastructure (NUFI) through the ESS-Lighthouse Q-MAT.
@@ A.1 Primitive lattice
$M_{\mathrm{prim}}\equiv\begin{pmatrix}\vec{m}_{1}&\vec{m}_{2}&\vec{m}_{3}\end{pmatrix}^{T}=\begin{pmatrix}0&1/a&1/a\\ 1/b&0&1/b\\ 1/c&1/c&0\end{pmatrix}^{T}.$ (22)
@@
where $\mathrm{trig}^{n}k=\cos k,\sin k,-\cos k,-\sin k$ for $n=0,1,2,3$ respectively, and $n$ should be evaluated modulo 4. This shows that the normal-state Hamiltonian can be written in terms of functions of integer multiples of the primitive momentum coordinates only, which means that we can obtain the corresponding real-space tight-binding model simply by replacing each Fourier coefficient $\mathrm{e}^{-ik_{i}}$ with its corresponding real-space translation operator $T_{i}$.
@@ A.2 Transformation to (0-11) coordinate system
$\hat{n}=\frac{-c\hat{e}_{y}+b\hat{e}_{z}}{\sqrt{b^{2}+c^{2}}},$ (26)
where we use $\vec{a}=\hat{e}_{x}a$, $\vec{b}=\hat{e}_{y}b$ and $\vec{c}=\hat{e}_{z}c$. Here, $a=4.161\>\mathrm{Å}$, $b=6.122\>\mathrm{Å}$ and $c=13.955\>\mathrm{Å}$, meaning that the angle between $\vec{c}$ and the vector normal to the surface is given by $\cos\theta=c/\sqrt{b^{2}+c^{2}}$ so $\theta=23.7^{\circ}$.
To properly capture the physics at the (0-11) surface, we first need to transform to a coordinate system that respects the symmetry of the surface, that is, we need two lattice vectors that are orthogonal to $\hat{n}$ and an integer multiple of the primitive lattice vectors. It is convenient to choose these as $\vec{r}_{1}$ and $\vec{r}_{+}=\vec{r}_{2}+\vec{r}_{3}$. We choose the last lattice vector of the coordinate system as $\vec{r}_{-}=\vec{r}_{3}-\vec{r}_{2}=\vec{b}-\vec{c}$, defining the full (0-11) coordinate system as
$R_{0-11}=\begin{pmatrix}\vec{r}_{1}&\vec{r}_{-}&\vec{r}_{+}\end{pmatrix}=\begin{pmatrix}-a/2&0&a\\ b/2&b&0\\ c/2&-c&0\end{pmatrix},$ (27)
with the corresponding reciprocal lattice vectors
$M_{0-11}=\begin{pmatrix}\vec{m}_{1}&\vec{m}_{-}&\vec{m}_{+}\end{pmatrix}=\begin{pmatrix}0&0&1/a\\ 1/b&1/2b&1/2b\\ 1/c&-1/2c&1/2c\end{pmatrix}.$ (28)
The Hamiltonian written in this coordinate system is $2\pi$ periodic and can thus be written as a function of $\exp{-ik_{i}}$ ) for $k_{i}=k_{1},k_{-},k_{+}$. Therefore, it corresponds to a real-space model with the mapping $\exp{-ik_{i}}\mapsto T_{i}$ where $T_{i}$ acts by translating with the lattice vector $\vec{r}_{i}=\vec{r}_{1},\vec{r}_{-},\vec{r}_{+}$ as
$T_{i}\ket{\vec{r}}=\ket{\vec{r}+\vec{r}_{i}}.$ (29)
However, since there is no integer combination of the new set of lattice vectors that leads to lattice transformations $\vec{c}$ or $\vec{b}$, we need to use a larger unit cell with another sublattice degree of freedom to properly represent the corresponding lattice translation operators $T_{y/z}$. The BdG Hamiltonian in the (0-11) coordinate system is therefore written in the basis
|$\ket{\vec{r}=n_{1}\vec{r}_{1}+n_{-}\vec{r}_{-}+n_{+}\vec{r}_{+}}\otimes\ket{\rho=p/h}\otimes\ket{\alpha=A/B}\otimes$
|$\ket{s=\mathrm{U}/\mathrm{Te}}\otimes\ket{\beta=0/1}\otimes\ket{\sigma=\uparrow/\downarrow},$ (30)
where |$\ket{\vec{r}=n_{1}\vec{r}_{1}+n_{-}\vec{r}_{-}+n_{+}\vec{r}_{+}}$ is the real-space basis vector, |$\ket{\rho=p/h}$ is the particle/hole sector from the BdG description, |$\ket{\alpha=A/B}$ is the sublattice from the (0-11) coordinate system transformation, |$\ket{s=\mathrm{U}/\mathrm{Te}}$ are the uranium/tellurium sectors, |$\ket{\beta=0/1}$ is the original sublattice, and |$\ket{\sigma=\uparrow/\downarrow}$ represents the pseudospin of the electrons.
We can now perform a Fourier transformation along the parallel momenta to obtain the Hamiltonian in the form
$H(\mathbf{k}^{\parallel})=H_{0}(\mathbf{k}^{\parallel})+T_{-}V(\mathbf{k}^{\parallel})+T_{-}^{T}V^{\dagger}(\mathbf{k}^{\parallel}).$ (31)
With the Hamiltonian written in the form from Eq. (31) we can use the method described in Ref. [47] to obtain the Green’s function components on the surface and bulk of the system with respect to the $\vec{r}_{-}$ direction. That is, we obtain a triple of Green’s functions $(G_{s}(\omega,\mathbf{k}^{\parallel}),G_{b}(\omega,\mathbf{k}^{\parallel}),\tilde{G}_{{s}}(\omega,\mathbf{k}^{\parallel}))$ as a function of energy $\omega$ and parallel momentum
$\mathbf{k}^{\parallel}=k_{1}\vec{m}_{1}+k_{+}\vec{m}_{+}.$ (32)
@@
where $m_{a}^{*}$ and $\vec{m}_{c}^{*}$ are orthogonal. Since $\vec{m}_{a}^{*}=\hat{e}_{x}/a$ we identify $k_{a}^{*}=k_{x}$. In the coordinates defined by Eq. (33), the reciprocal lattice vectors are given by $(2\pi,0)$ and $(\pi,2\pi)$.
@@
$\mathcal{H}=\bigoplus_{n=0}^{N}\mathcal{H}_{n},$ (34)
with basis vectors |$\ket{n,\alpha}$ where $\alpha$ labels a complete basis on the $\mathcal{H}_{n}$ subspaces. The Hamiltonian is represented as
@@
Note that this system exhibits translational invariance except at the boundaries of the system. Note also that it is possible to bring any system that has translation symmetry except at the boundary into this form simply by extending the size of the unit cell until each cell is connected only through nearest-neighbor interactions. We now wish to derive an iterative procedure to find the Green’s function in the first block of the system, as described in Ref. [47].
@@
$P_{n}=\sum_{\alpha}\outerproduct{n,\alpha}{n,\alpha}.$ | . (40)
@@
$\alpha_{i}=\alpha_{i-1}(\omega-\varepsilon_{i-1})^{-1}\alpha_{i-1},$ (48)
$\beta_{i}=\beta_{i-1}(\omega-\varepsilon_{i-1})^{-1}\beta_{i-1},$ (49)
$\varepsilon_{i}=\varepsilon_{i-1}+\alpha_{i-1}(\omega-\varepsilon_{i-1})^{-1}\beta_{i-1}+\beta_{i-1}(\omega-\varepsilon_{i-1})^{-1}\alpha_{i-1},$ (50)
$\varepsilon_{i}^{s}=\varepsilon_{i-1}^{s}+\alpha_{i-1}(\omega-\varepsilon_{i-1})^{-1}\beta_{i-1},$ (51)

### v2

@@ III.2 Superconducting topological surface states
As a result of the odd parity of the superconducting order parameter, the boundaries of this topological system may exhibit zero-energy bound states in the form of Majorana-Dirac surface cones, that is, states with linear dispersion that form around time-reversal invariant momenta (TRIM) in the surface BZ [48, 2, 49, 25, 50, 28, 27].
@@
Figure 6: Spectral functions for the B2u (a) and B3u (b) phases while moving through the BZ path $\GammaXSYΓ\GammaS$ as defined in Fig. 7(b). Bulk bands are displayed in red (uranium) and blue (tellurium) colors, highlighting the nature of the nodes on the surface BZ. The green bands show the topological surface states on the (0-11) surface.
In order to gain insight into the topological nature of the surface bound states and determine whether there are any qualitative differences between the B2u versus B3u phases, we consider the symmetry representations for the two different phases. We will exploit that the (0-11) surface preserves only the $\mathcal{M}_{x}$ crystal symmetry along with time-reversal $\mathcal{T}$ and the particle-hole anti-symmetry $\mathcal{P}$. Since the characters of the B2u and B3u phases are opposite under $\mathcal{M}_{x}$, this leads to qualitative differences that are robust against microscopic changes of the model that preserve $\mathcal{M}_{x}$. We note that the even/odd property of B3u/B2u under $\mathcal{M}_{x}$ implies that they do not mix on the (0-11) surface. Their distinct symmetries under $\mathcal{M}_{x}$ can be seen directly from Fig. 4 with B2u (B3u) being odd (even) under $\mathcal{M}_{x}$. Specifically, we will see that the B2u phase features Majorana flat bands around the $\Gamma$ point, while the B3u exhibits a Majorana-Dirac surface cone with linear dispersion in both $k_{x}$ and $k_{c^{*}}$.
First, we consider the algebra of the representations of the symmetries in the bulk. Following Ref. 50, we express the commutation relation between the BdG representation $U$ of the crystal symmetry $g$ and the particle-hole anti-symmetry $\mathcal{P}$ in terms of the character $\chi(g)$ as
$U(g)U(\mathcal{P})=\chi(g)U(\mathcal{P})\mathcal{U}(g),\$(17)
where the representation of particle-hole is given by $U(\mathcal{P})=\tau_{1}K$, where $\tau_{1}$ is the first Pauli matrix in particle-hole space and $K$ is complex conjugation. Due to the presence of strong SOC, the spin and momenta transform together, resulting in the normal state representation of the $\mathcal{M}_{x}$ symmetry given by $i\sigma_{x}$, where $\sigma_{i}$ are the Pauli matrices in spin-space. Both the bulk and the surface respect time-reversal symmetry and particle-hole anti-symmetry and therefore also the chiral anti-symmetry defined as $\mathcal{C}=\mathcal{T}\mathcal{P}$ [51]. Note that chiral symmetry leaves the momentum invariant, acting only on the internal degrees of freedom. Since time-reversal symmetry $U(\mathcal{T})=i\sigma_{y}K$ commutes with $\mathcal{M}_{x}$, the commutation relation between $\mathcal{C}$ and $\mathcal{M}_{x}$ are given by
@@
where $\chi(\mathcal{M}_{x})=1$ for B2u and $\chi(\mathcal{M}_{x})=-1$ for B3u as seen from Tab. 6. In the basis of the Majorana zero modes, a similar relation must hold for the representation of the symmetries of the surface, which in turn constrains the dispersion of the surface states. To see this, we write an effective surface Hamiltonian $H_{s}(\mathbf{k^{\parallel}})$ around the $\Gamma$ point. In both the B2u and B3u cases, we observe two localized surface states and we can therefore write it as a two-band model in terms of the Pauli matrices $\eta_{i}$. We denote the symmetry representation in the basis of the surface states as $\mathcal{O}(g)$, where $g=\mathcal{M}_{x},\mathcal{P},\mathcal{T}$ are the set of symmetry operations preserved by the (0-11) cleave plane. We define the basis such that time-reversal and particle-hole representations are given by
$\mathcal{O}(\mathcal{P})=\eta_{0}K,\$
@@
The above argument relies on the fact that only one crystal symmetry remains on the surface, under which B3u and B2u transform differently. This should be contrasted to a related discussion in Ref. 28 where the authors calculate the surface states on the (001) surface. That is, a surface preserving both $\mathcal{M}_{x}$ and $\mathcal{M}_{y}$ symmetries, such that no qualitative difference can be made between the B2u and B3u phases. Indeed, they find that both phases generically host flat bands around the $\Gamma$ point [28].
@@ IV Discussion and Conclusions
We note that a recent experimental STM quasiparticle interference (QPI) study reported enhanced scattering intensity in the superconducting phase at specific scattering wavevectors [38]. In an accompanying work, we have developed the theoretical framework for addressing QPI in spin-triplet superconductors and combined it with the current model for superconductivity in UTe2 on the (0-11) surface [56]. A main conclusion of that study is that B3u appears more consistent with the currently available QPI data. We refer to Ref. 56 for further details of the theoretical QPI analysis relevant for UTe2.
@@ Acknowledgements.
We acknowledge fruitful discussions with Daniel Agterberg, J. C. Séamus Davis, Peter J. Hirschfeld, and Shuqui Wang. H. C. acknowledges support from the Novo Nordisk Foundation grant NNF20OC0060019. A.K. acknowledges support by the Danish National Committee for Research Infrastructure (NUFI) through the ESS-Lighthouse Q-MAT.
@@ A.1 Primitive lattice
$M_{\mathrm{prim}}\equiv\begin{pmatrix}\vec{m}_{1}&\vec{m}_{2}&\vec{m}_{3}\end{pmatrix}^{T}=\begin{pmatrix}0&1/a&1/a\\ 1/b&0&1/b\\ 1/c&1/c&0\end{pmatrix}^{T}.\$(22)
@@
where $\mathrm{trig}^{n}k=\cos k,\sin k,-\cos k,-\sin k$ for $n=0,1,2,3$ respectively, and $n$ should be evaluated modulo 4. This shows that the normal-state Hamiltonian can be written in terms of functions of integer multiples of the primitive momentum coordinates only, which means that we can obtain the corresponding real-space tight-binding model simply by replacing each Fourier coefficient $\mathrm{e}^{-ik_{i}}$ with its corresponding real-space translation operator $T_{\vec{r}_{i}}$.
@@ A.2 Transformation to (0-11) coordinate system
[…]
where we use $\vec{a}=a\hat{e}_{x}$, $\vec{b}=b\hat{e}_{y}$ and $\vec{c}=c\hat{e}_{z}$ and $a=4.161\>\mathrm{Å}$, $b=6.122\>\mathrm{Å}$ and $c=13.955\>\mathrm{Å}$ are the lattice constant along the three crystal axes [53]. The angle between $\hat{e}_{y}$ and the vector normal to the surface is given by $\cos\theta=c/\sqrt{b^{2}+c^{2}}$ with $\theta=23.7^{\circ}$.
To properly capture the physics at the (0-11) surface, we first need to transform to a coordinate system that respects the symmetry of the surface, that is, we need two lattice vectors that are orthogonal to $\hat{n}$ and an integer multiple of the primitive lattice vectors. It is convenient to choose these as $\vec{r}_{1}$ and $\vec{r}_{+}=\vec{r}_{2}+\vec{r}_{3}$ and we choose the last vector of the coordinate system as $\vec{r}_{3}$. We can now perform a Fourier transformation in the parallel momenta to obtain a Hamiltonian in the form
$H(\mathbf{k}^{\parallel})=H_{0}(\mathbf{k}^{\parallel})+T_{\vec{r}_{3}}V(\mathbf{k}^{\parallel})+T_{\vec{r}_{3}}^{T}V^{\dagger}(\mathbf{k}^{\parallel}),\$(27)
where $T_{\vec{r}_{3}}$ is the real space operator that translates by $\vec{r}_{3}$. From this form, we can use the method described in Ref. 47 to obtain the Green’s function components on the surface and bulk of the system with respect to the $\vec{r}_{3}$ direction. That is, we obtain a triple of Green’s functions $(G_{s}(\omega,\mathbf{k}^{\parallel}),G_{b}(\omega,\mathbf{k}^{\parallel}),\tilde{G}_{{s}}(\omega,\mathbf{k}^{\parallel}))$ as a function of energy $\omega$ and parallel momentum
$\mathbf{k}^{\parallel}=k_{1}\vec{m}_{1}+k_{+}\vec{m}_{+}.\$(28)
@@
where $\vec{m}_{a^{*}}$ and $\vec{m}_{c^{*}}$ are orthogonal. We identify $k_{a^{*}}=k_{x}$ since $\vec{m}_{a^{*}}=\hat{e}_{x}/a$ while $\vec{m}_{c}^{*}=\hat{e}_{y}/b+\hat{e}_{z}/c$. In the coordinates defined by Eq. (29), the reciprocal lattice vectors are given by $(2\pi,0)$ and $(\pi,2\pi)$.
Note that in the figures in the main text, we reinstate the units of the momentum vectors, such that $k_{x}$ is measured in units of $1/a$ and $k_{c^{*}}$ in units of $1/c^{*}$ with $c^{*}=\sqrt{c^{2}+b^{2}}/2$·
@@
$\mathcal{H}=\bigoplus_{n=0}^{N}\mathcal{H}_{n},\$(30)
[…]
@@
Note that this system exhibits translational invariance except at the boundaries of the system. Note also that it is possible to bring any system that has translation symmetry except at the boundary into this form simply by extending the size of the unit cell until each cell is connected only through nearest-neighbor interactions. We now wish to derive an iterative procedure to find the Green’s function in the first block of the system, as described in Ref. 47.
@@
$|.P_{n}=\sum_{\alpha}\outerproduct{n,\alpha}{n,\alpha}$.\ (36)
@@
$\alpha_{i}=\alpha_{i-1}(\omega-\varepsilon_{i-1})^{-1}\alpha_{i-1},\$(44)
$\beta_{i}=\beta_{i-1}(\omega-\varepsilon_{i-1})^{-1}\beta_{i-1},\$(45)
$\varepsilon_{i}=\varepsilon_{i-1}+\alpha_{i-1}(\omega-\varepsilon_{i-1})^{-1}\beta_{i-1}+\beta_{i-1}(\omega-\varepsilon_{i-1})^{-1}\alpha_{i-1},\$(46)
$\varepsilon_{i}^{s}=\varepsilon_{i-1}^{s}+\alpha_{i-1}(\omega-\varepsilon_{i-1})^{-1}\beta_{i-1},\$(47)